6. Calculating simple interest and APR on a single-payment loan Aa Aa E You are taking out a single-payment loan that uses the simple interest method to compute the finance charge. You need to figure out what your payment will be when the loan comes due. The equation to calculate the finance charge is: In the equation, Fs is the finance charge for the loan. What are the other values? P is the r is the stated t is the term of the loan in amount of the loan. rate of interest. Youre borrowing $10,000 for a year with a stated annual interest rate of 6%. Complete the following table. (Note: Round your answers to the nearest dollar.) Principal Finance charge D Total Payback D $10,000 Annual Percentage Rate (APR) You also want to calculate the APR (annual percentage rate) and compare it to the stated interest rate. Average Annual Finance Charge Average Loan Balance Outstanding APR = First, compute the average annual finance charge by dividing the total finance charge of $ the loan, which is a year (1.0 year) s by the life of (Note: Round your answers to the nearest dollar).

Assignment Chapter 07- Using Consumer Loans Annual Percentage Rate (APR) You also want to calculate the APR (annual percentage rate) and compare it to the stated interest rate. Average Annual Finance Charge Average Loan Balance Outstanding APR First, compute the average annual finance charge by dividing the total finance charge of s by the life of (Note: Round your answers to the nearest dollar). Next, as a single-payment loan, the average loan balance outstanding is constant at s10,000. Complete the calculation. (Note: Round your answers to the nearest dollar or whole perc entage point.) Average Annual Finance Charge Average Loan Balance outstanding because the: The stated interest rate and APR are O Simple interest method was used to calculate finance charges O Loan is a single-payment loan O Term of the loan is fewer than five years
Answer #1

Fs=P*r*t

P is the principal amount of the loan

r is the stated annual rate of interest

t is the term of the loan in years

Finance charge=10000*6%=600

Total payback=10600

Dividing the total finance charge of $600 by the life of the loan which is a year=$600

APR=600/10000=6%

Loan is a single payment loan